# Pixels to Prices: Visual Traits, Market Cycles, and the Economics of NFT Valuation

### v1

@@ Abstract
This paper studies how visual traits and market cycles shape prices in NFT markets. Using 94,039 transactions from 26 major generative Ethereum collections, the analysis extracts 196 machine-quantified image features—covering color, composition, palette structure, geometry, texture, and deep-learning embeddings—then applies a three-stage filter process to identify stable predictors for hedonic regression. A static mixed-effects model shows that market sentiment and transparent, interpretable image traits have significant and independent pricing power: higher focal saturation, compositional concentration, and curvature are rewarded, while clutter, heavy line work, and dispersed palettes are discounted; deep embeddings add limited incremental value conditional on explicit traits. To assess state dependence, the study estimates a Bayesian dynamic mixed-effects panel with cycle effects and time-varying coefficients for a salient image attribute (Composition Focus—Saturation). Collection-level heterogeneity (“brand premia”) is absorbed by random effects. The time-varying coefficients exhibit regime sensitivity, with stronger premia in expansionary phases and weaker or negative loadings in downturns, while grand-mean effects remain small on average. Overall, NFT prices reflect both observable digital product characteristics and market regimes, and the framework offers a cycle-aware tool for asset pricing, platform strategy, and market design in digital art markets.
@@ 3.1 NFT Collections and Transaction Data
3.2 Image Data and Feature Extraction
A large-scale corpus of 24,148 NFT images spans the study window. Each image is mapped to a unified feature table combining interpretable “classic” vision descriptors with compact deep embeddings.
@@ 3.2 Image Data and Feature Extraction
Concretely, the classic block covers (i) color statistics in HLS/HSV space (including dominant–hue diagnostics and contrast ranges); (ii) composition–focus ratios comparing the central region to the full canvas (hue, lightness, saturation); (iii) edge geometry and strength (Canny/Sobel coverage, horizontal/vertical span); (iv) palette structure from a 6–color K–means (e.g., saturation/value moments, Lab colorfulness, mean $\Delta E$); (v) line–art structure (straight vs. curved prevalence, fractal dimension, stroke thickness proxies); and (vi) two texture families: a 10–bin Local Binary Pattern (LBP) histogram and five radial frequency–band energies from the image FFT.
To capture global content and style, deep representations are included: a 2048–D ResNet descriptor reduced to 100 principal components, and Gram–matrix style features from VGG layers reduced to 50 components. The resulting inventory totals 196 candidate features per image prior to modeling. [^1] Implementation details include batched inference for CNNs, ImageNet normalization, and consistent pre-processing (resize to 256, center–crop 224). A compact, collection–agnostic script outputs a single CSV (row = image) with all classic, texture, and deep/PCA features.
Two practical choices are worth noting. First, hue is circular; whenever hue enters a regression it is encoded via $\sin/\cos$ and the raw angle is omitted to avoid discontinuity and collinearity. Second, no dataset–level standardization is performed inside the feature script; for econometric stability, all numeric regressors are z–scored after merging with returns/sentiment data in the analysis pipeline.
Not all 196 features enter the final regressions. To safeguard parsimony and out–of–sample stability, a small, pre–registered reduction step screens highly collinear candidates and retains a compact subset for the static and dynamic hedonic models. In short: the data work is heavy, the modeling is disciplined. Extracting a hundred–plus descriptors per image across tens of thousands of tokens enables richer questions about how visible design choices and style map into prices—without overfitting the noise.

### v2

@@ Abstract
Pixels and market cycles both move NFT prices. Using 94,039 transactions from 26 major generative Ethereum collections, this study extracts 196 machine-quantified image descriptors—color, composition, palette structure, geometry, texture, and deep-learning embeddings—and applies a three-stage filter to identify stable predictors for hedonic regression. A static mixed-effects model shows that market sentiment and transparent, interpretable image traits have significant and independent pricing power: higher focal saturation, tighter compositional concentration, and greater curvature are rewarded, while clutter, heavy line work, and dispersed palettes are discounted; deep embeddings add limited incremental value once explicit traits are included. To assess state dependence, a Bayesian dynamic mixed-effects panel with cycle effects is estimated, allowing Composition Focus—Saturation—the ratio of saturation in the central region to the whole image, capturing vividness and concentration at the focal area—to vary across market regimes. Collection-level heterogeneity (brand premia) is absorbed by random effects. The time-varying coefficients exhibit clear regime sensitivity, with stronger premia in expansionary phases and weaker or negative loadings in downturns, while the grand-mean effect is small on average. Overall, NFT prices reflect both observable digital product characteristics and market regimes, and the framework offers a cycle-aware tool for asset pricing, platform strategy, and market design in digital art markets.
@@ 3.1 NFT Collections and Transaction Data
3.2 Image Feature Extraction
Each image is encoded as a 196–dimensional vector comprising three complementary feature families: 31 classic, interpretable cues; 15 texture descriptors; and 150 compact deep representations. The aim is to summarize visible design choices in a way that is both interpretable and model–ready; brief “for example” explanations follow each technical item.
Classic features (31 total). Concretely, the classic block covers: (i) color statistics in HLS/HSV (8), including global hue, lightness, and saturation with diversity diagnostics such as hue range, number of distinct hue bins, and the modal hue with its share. For example: a bright neon background exhibits high saturation; a washed-out pastel exhibits low saturation; a predominantly golden figure has a hue near yellow/orange. (ii) composition–focus ratios (3) comparing the central third of the canvas to the full image for hue, lightness, and saturation (COMPOSITION_FOCUS_*); values >1>1 indicate concentrated focal design, while <1<1 indicates diffuse layout. For example: a centered character on a plain background yields a ratio >1>1, whereas a uniform gradient yields $\approx 1$. (iii) edge geometry and strength (5) via Canny coverage, horizontal/vertical span (EDGE_RANGE_X/Y), bounding-box share, and a Sobel energy ratio (edge salience relative to the image mean). For example: thick inked outlines raise edge area and energy; soft airbrushed shading lowers both. (iv) palette structure from a 6–color K–means (6), summarized by HSV saturation/value moments, Lab "colorfulness," and mean $\Delta E$ separation among palette centroids. For example: cohesive pastels show small $\Delta E$; clashing neons show large $\Delta E$.
@@ 3.2 Image Feature Extraction
(v) line-art geometry (6) using skeletonization and Hough segments to proxy average stroke length and thickness, straight vs. curved prevalence, a box-counting fractal dimension, and endpoints per segment. For example: ruler-straight hatching increases straight-segment prevalence; organic drawings increase curved structure. (vi) single-number “quality” scalars (3): Laplacian variance (sharpness), Shannon entropy (tonal richness), and histogram energy (concentration). For example: crisp, high-detail drawings have higher sharpness; nearly flat backgrounds have lower entropy but higher energy (mass concentrated in a few gray levels).
Texture features (15 total). Two complementary texture families are used: Local Binary Patterns (10-bin histogram) summarize micro-textures; and FFT ring energies (5 concentric frequency bands) split spectral power from coarse to fine detail. For example: a polka-dot hoodie raises specific LBP bins; fine stippling increases high-frequency FFT bands, whereas broad flat fills increase low-frequency bands.
Deep representations (150 total). To capture global content and style, deep representations are included: a 2048–D ResNet descriptor reduced to 100 principal components, and Gram–matrix style features from VGG layers reduced to 50 components. The resulting inventory totals 196 candidate features per image prior to modeling. [^1] Implementation details include batched inference for CNNs, ImageNet normalization, and consistent pre–processing (resize to 256, center–crop 224). A compact, collection–agnostic script outputs a single CSV (one row per image) with all classic, texture, and deep/PCA features.
To summarize, classic features align with recognizable design choices (focal saturation, edge salience, curvature, palette cohesion), allowing coefficients to be read as premia/discounts on these attributes, while deep PCs act as orthogonal controls that absorb remaining visual regularities without over-parameterizing the regression.
Two practical choices are worth noting. First, hue is circular; whenever hue enters a regression it is encoded via $\sin/\cos$ and the raw angle is omitted to avoid discontinuity and collinearity. Second, no dataset-level standardization is performed inside the feature script; for econometric stability, all numeric regressors are z-scored after merging with returns/sentiment data in the analysis pipeline.
Not all 196 features enter the final regressions. To safeguard parsimony and out-of-sample stability, a small, pre-registered reduction step screens highly collinear candidates and retains a compact subset for the static and dynamic hedonic models. Extracting a hundred-plus descriptors per image across tens of thousands of tokens enables richer questions about how visible design choices and style map into prices—without overfitting the noise.
@@ 5.2 Dynamic hedonic model with time-varying parameters
Figure 2: Time-varying premium for Composition Focus—Saturation from the Bayesian TVP model. Points are posterior means of the semi-elasticity of $\log(1+\text{price})$ per one–standard-deviation change in the standardized regressor; bands are 50% and 95% credible intervals. The path is near zero on average, modestly positive around the Post Blur Lull 2023 state, and slightly negative in Mid 2024 Ebb and Q1 2025 Slump, consistent with state-dependent pricing.
[…]
@@ Appendix A. Additional Tables
Table A3: Time-varying states for Composition Focus Saturation
cycle mean
ATH_Blowoff_Sep21_Jan22 -0.004
Blur_Surge_Feb_Mar_2023 0.012
Crypto_Winter_Terra_FTX -0.005
Early_Boom_2021H1 -0.049
Mid_2024_Ebb -0.008
Post_Blur_Lull_2023 0.018
Pre_Crash_Softening_2022Q1 0.001
Q1_2025_Slump -0.008
Q4_2024_Rebound 0.011
Thaw_Q4_2023_Q1_2024 -0.006
Table A3: Time-varying states for COMPOSITION_FOCUS_SATURATION (full listing).
Table A4: Image Feature Categories and Descriptions (Part 1: Classic Features)
Feature Category Count Description and Examples
Color in HLS/HSV 8 Captures global hue, lightness, and saturation alongside diversity statistics (hue range, number of distinct hue bins, modal hue and its share).
Example: Bright neon background has high saturation; washed-out pastel has low saturation; predominantly golden figure has hue near yellow/orange.
Composition focus: center vs. whole 3 Uses ratios comparing central third of canvas to full image for hue/lightness/saturation (COMPOSITION_FOCUS_*). Values >1>1 indicate concentrated focal area, <1<1 indicates diffuse layout.
Example: Centered character on plain background yields ratio >1>1; uniform gradient background yields $\approx 1$.
Edges via Canny/Sobel 5 Measures fraction of edge pixels and their spatial span (bounding-box share, horizontal/vertical range), together with Sobel energy ratio (edge strength relative to image mean).
Example: Thick inked outlines produce higher edge area and stronger energy; soft airbrushed shading produces lower values.
Single-number "quality" scalars 3 Includes Laplacian variance (sharpness), Shannon entropy (tonal richness), and histogram energy (concentration) for complementary one-number summaries.
Example: Crisp, high-detail drawing has higher sharpness; nearly flat background has lower entropy but higher energy (most mass in few gray levels).
Palette structure with K=6K{=}6 colors 6 Applies K-means on RGB to isolate six dominant colors; from their HSV/Lab centroids, computes saturation/value moments, Lab "colorfulness," and mean $\Delta E$ separation.
Example: Cohesive pastel set shows small $\Delta E$; mix of clashing neons shows large $\Delta E$.
Line-art geometry 6 Uses skeletonization and Hough segments to proxy average stroke length and thickness, straight vs. curved prevalence, box-counting fractal dimension, and endpoints per segment.
Example: Technical, ruler-straight hatching yields more straight segments; organic drawings yield more curved structure and different fractal score.
Classic Features Total: 31 per image
  • Notes: All features are extracted from 24,148 NFT images across 26 collections. Classic features provide interpretable design choices and are standardized (z-scored) before entering regression models.
Table A5: Image Feature Categories and Descriptions (Part 2: Texture and Deep Features)
Feature Category Count Description and Examples
Texture: LBP histogram and FFT ring energies 15 Employs Local Binary Patterns to summarize micro-textures (10-bin histogram); concentric FFT bands split spectral power from coarse to fine detail.
Example: Polka-dot hoodie pushes specific LBP bins up; fine stippling increases high-frequency FFT bands, whereas broad flat fills increase low-frequency bands.
Deep representations for global content and style 150 Utilizes pre-trained CNNs with standardized pre-processing (resize to 256, center-crop 224, ImageNet normalization). Global 2048-dimensional descriptor from ResNet-50 (average-pooled convolutional activations) reduced by PCA to 100 components (CNN_PCA_001–100), summarizing object/shape/layout information; in parallel, VGG-19 Gram-matrix features (second-order co-activations within selected convolutional blocks) reduced to 50 components (STYLE_GRAM_PCA_001–050), emphasizing style/texture and color co-occurrence.
Example: Content PC may capture "frontal face vs. profile" or "figure vs. background layout," while style PC separates "flat vector look" from "brushy/stippled texture."
Total Features: 196 per image (31 classic + 15 texture + 150 deep representations)
  • Notes: Deep representations capture complex visual patterns that complement the interpretable classic features. All features are standardized (z-scored) before entering regression models.